# Machine Learning and the Physical Sciences

### v1

@@ 1 Introduction
Galaxy morphology concepts are traditionally encoded in manually-designed taxonomies, e.g., DeVaucouleurs1963 . These may miss concepts that are too rare to be manually detected. New, rare concepts are statistically guaranteed to be represented in the billions of resolved galaxies soon to be imaged by space telescopes like Euclid and Roman (scaramellaEuclidPreparationEuclid2021, ; spergel2015widefieldinfrarredsurveytelescopeastrophysics, ), so we will need a data-driven method to identify these rare concepts. Relatedly, the scale of Euclid and Roman also demands an increasing reliance on deep learning models for interpreting images (e.g. classification, segmentation, feature extraction for multimodal tasks, etc.) Identifying the concepts learned by our models can inform how we build them and help mitigate the risk of unintended biases or shortcut learning (geirhosShortcutLearningDeep2020, ).
Sparse autoencoders (SAEs) are routinely used for interpreting language models by identifying the set of directions in activation space that can together describe any activation vector in a neural network (makhzani2014ksparseautoencoders, ). Sparsity encourages these directions to be monosemantic, i.e., corresponding to a single concept (elhage2022toymodelssuperposition, ; cunninghamSparseAutoencodersFind2023, ; 2025ApJ…980..183W, ; karvonenSAEBenchComprehensiveBenchmark2025, ). Here, we apply SAEs to reveal the concepts learned by galaxy morphology models.
We construct embeddings for galaxies in the Euclid Q1 data release (collaborationEuclidQuickData2025, ) and then compress those embeddings via SAEs and PCA to identify learned concepts. We show that SAEs efficiently extract candidate morphological features from pretrained networks, comparing against a PCA baseline. For both supervised and self-supervised embeddings, SAEs identify interpretable features aligned with, and beyond the Galaxy Zoo decision trees, while PCA primarily aligns with existing categories.
@@ 2 Representations of Galaxy Morphology
Euclid Imaging and Embeddings. We analyze two types of neural network embeddings for galaxy images in the Euclid Q1 dataset. For supervised embeddings, we use the Zoobot models presented in euclidcollaborationEuclidQuickData2025b . These models are pretrained on approximately 1M volunteer-annotated galaxies from four other telescopes and then finetuned using 170k Euclid galaxies with new annotations from Galaxy Zoo Euclid (GZ). Our chosen model, ConvNeXT-Nano, has an encoder with 640-dimensional features.
For self-supervised embeddings, we train a masked autoencoder (MAE, (heMaskedAutoencodersAre2022, )) on 3M Euclid images drawn from the internal Euclid dataset “RR2”. This is $5-10\times$ larger than recent work applying MAE to astronomy images (fathkouhiAstroMAERedshiftPrediction2024, ) and applying other self-supervised learning approaches to Euclid images (euclidcollaborationEuclidQuickData2025, ).
@@
We use a high masking fraction of 90% (vs. 75% as published). We find that networks of just 1M parameters perform well with conventional masking ratios, suggesting that astronomy images have high redundancy. Work using MAE for video (which is intrinsically high redundancy) shows that higher masking ratios are required to force the network to do more than interpolate (feichtenhoferMaskedAutoencodersSpatiotemporal2022, ).
@@
We adjust the frequency of the positional encoding to indicate the angular scale of each image. This is an astronomical analogy to work using MAE for remote sensing (reedScaleMAEScaleAwareMasked2023, ), who adjust the frequency to indicate the ground scale of each satellite image.
Our trained MAE reconstructs masked galaxy features more accurately than professional astronomers. Figure 1 shows our MAE infilling 90% masked galaxy images with near-perfect accuracy. An informal survey of our human colleagues suggests that professional astronomers cannot correctly describe the same masked images. Superhuman reconstruction performance indicates that our MAE extracts meaningful features. We also hope that our MAE will be useful for tasks like patch-level anomaly search (find me images that include patches like these) and targeted infill (remove this small artifact). We release our MAE model on HuggingFace at https://huggingface.co/mwalmsley/models.
@@
Extracting Features from Embeddings. SAEs learn sparse representations of the embeddings using an overcomplete basis. We train a Matryoshka Sparse Autoencoder (SAE; (bussmann2025learningmultilevelfeaturesmatryoshka, )) with hierarchical group sizes 64, 64, 128, $\cdots$, 1024, batch top-$k$ sparsity ($k=64$), a small L1 penalty, and auxiliary loss for reviving dead neurons (following (2024arXiv240800657O, )). For both self-supervised and supervised embeddings, we train for 200 epochs. The code is publicly available at https://github.com/jwuphysics/euclid-galaxy-morphology-saes.
@@ 3 Results
Self-supervised: The MAE embeddings exhibit surprisingly strong correlations with GZ—they are more correlated than the Zoobot embeddings. We suspect this arises because the MAE embeddings are intent on reconstructing imaging artifacts such as saturated stars or ghosts, which can be diverse and span many pixels. The SAE and PCA features primarily activate on these artifacts, and indeed we see that the most of the strongly aligned GZ classes are problem_* or artifact_*. [^3] Another possibility is that the supervised optimization objective (i.e. aligning to GZ) can (counterintuitively) cram class-separating features into a low-dimensional subspace (e.g., grigg2021selfsupervisedsupervisedmethodslearn, ). In other words, the supervised objective can cause embeddings to span a small number of known, dominant morphological classes, whereas the self-supervised case allows representations to disentangle and occupy a broader space of potentially novel features.
@@ 4 Discussion
Limitations. Unfortunately, some SAE features remain difficult to interpret, despite the benefits of using Matryoshka SAEs to combat feature splitting and absorption (Chanin2024AIF, ; bussmann2025learningmultilevelfeaturesmatryoshka, ). Still, SAEs surpass our PCA baseline approach. PCA features become unidentifiable after $\sim 30$ features and show no alignment with GZ. In contrast, our SAE delivers an order of magnitude more aligned features (see Appendix A for more details).
Scientific Discovery. SAEs can extract interpretable features from both supervised and self-supervised models at scale (gao2024scalingevaluatingsparseautoencoders, ). Sparse activations are steerable and can be readily interpreted through automated means (cunninghamSparseAutoencodersFind2023, ; 2024arXiv240800657O, ; ye2024steering, ). Although there are concerns about whether SAEs are effective at learning known concepts, they excel at identifying new features compared to other methods (muhamed2024decodingdarkmatterspecialized, ; movva2025sparseautoencodershypothesisgeneration, ; marks2025auditinglanguagemodelshidden, ; peng2025usesparseautoencodersdiscover, ). Thus, SAEs have tremendous potential to aid scientific discovery by categorizing and sifting through next-generation astronomical datasets (see also, e.g., (2021A&C….3600481L, ; oryan2025identifyingastrophysicalanomalies996, )). By extracting features from self-supervised embedding models, SAEs can act as scalable discovery engines that surface rare or anomalous astrophysical phenomena for future targeted investigation.

### v2

@@ 1 Introduction
Galaxy morphology concepts are traditionally encoded in manually-designed taxonomies (e.g., De Vaucouleurs, 1963). These may miss concepts that are too rare to be manually detected. New, rare concepts are statistically guaranteed to be represented in the billions of resolved galaxies soon to be imaged by space telescopes like Euclid and Roman (Scaramella et al., 2021; Spergel et al., 2015), so we will need a data-driven method to identify these rare concepts. Relatedly, the scale of Euclid and Roman also demands an increasing reliance on deep learning models for interpreting images (e.g. classification, segmentation, feature extraction for multimodal tasks, etc.) Identifying the concepts learned by our models can inform how we build them and help mitigate the risk of unintended biases or shortcut learning (Geirhos et al., 2020).
Sparse autoencoders (SAEs) are routinely used for interpreting language models by identifying the set of directions in activation space that can together describe any activation vector in a neural network (Makhzani and Frey, 2014). Sparsity encourages these directions to be monosemantic, i.e., corresponding to a single concept (Elhage et al., 2022; Cunningham et al., 2023; Wu, 2025; Karvonen et al., 2025). Here, we apply SAEs to reveal the concepts learned by galaxy morphology models.
We construct embeddings for galaxies in the Euclid Q1 data release (Euclid Collaboration et al., 2025a) and then compress those embeddings via SAEs and PCA to identify learned concepts. We show that SAEs efficiently extract candidate morphological features from pretrained networks, comparing against a PCA baseline. For both supervised and self-supervised embeddings, SAEs identify interpretable features aligned with, and beyond the Galaxy Zoo decision trees, while PCA primarily aligns with existing categories.
@@ 2 Representations of Galaxy Morphology
Euclid Imaging and Embeddings. We analyze two types of neural network embeddings for galaxy images in the Euclid Q1 dataset. For supervised embeddings, we use the Zoobot models presented in Euclid Collaboration et al. (2025c). These models are pretrained on approximately 1M volunteer-annotated galaxies from four other telescopes and then finetuned using 170k Euclid galaxies with new annotations from Galaxy Zoo Euclid (GZ). Our chosen model, ConvNeXT-Nano, has an encoder with 640-dimensional features.
For self-supervised embeddings, we train a masked autoencoder (MAE, He et al. 2022) on 3M Euclid images drawn from the internal Euclid dataset “RR2”. This is $5-10\times$ larger than recent work applying MAE to astronomy images (Fathkouhi and Fox, 2024) and applying other self-supervised learning approaches to Euclid images (Euclid Collaboration et al., 2025b).
@@
We use a high masking fraction of 90% (vs. 75% as published). We find that networks of just 1M parameters perform well with conventional masking ratios, suggesting that astronomy images have high redundancy. Work using MAE for video (which is intrinsically high redundancy) shows that higher masking ratios are required to force the network to do more than interpolate (Feichtenhofer et al., 2022).
@@
We adjust the frequency of the positional encoding to indicate the angular scale of each image. This is an astronomical analogy to work using MAE for remote sensing (Reed et al., 2023), who adjust the frequency to indicate the ground scale of each satellite image.
Our trained MAE reconstructs masked galaxy features more accurately than professional astronomers. Figure 1 shows our MAE infilling 90% masked galaxy images with near-perfect accuracy. An informal survey of our human colleagues suggests that professional astronomers cannot correctly describe the same masked images. Superhuman reconstruction performance indicates that our MAE extracts meaningful features. We also hope that our MAE will be useful for tasks like patch-level anomaly search (find me images that include patches like these) and targeted infill (remove this small artifact). We release our MAE model on HuggingFace at https://huggingface.co/mwalmsley/euclid-rr2-mae. We also present a demo of the MAE at https://huggingface.co/spaces/mwalmsley/euclid_masked_autoencoder.
@@
Extracting Features from Embeddings. SAEs learn sparse representations of the embeddings using an overcomplete basis. We train a Matryoshka Sparse Autoencoder (SAE; (Bussmann et al., 2025)) with hierarchical group sizes 64, 64, 128, $\cdots$, 1024, batch top-$k$ sparsity ($k=64$), a small L1 penalty, and auxiliary loss for reviving dead neurons (following (O’Neill et al., 2024)). For both self-supervised and supervised embeddings, we train for 200 epochs. The code is publicly available at https://github.com/jwuphysics/euclid-galaxy-morphology-saes.
@@ 3 Results
Self-supervised: The MAE embeddings exhibit surprisingly strong correlations with GZ—they are more correlated than the Zoobot embeddings. We suspect this arises because the MAE embeddings are intent on reconstructing imaging artifacts such as saturated stars or ghosts, which can be diverse and span many pixels. The SAE and PCA features primarily activate on these artifacts, and indeed we see that the most of the strongly aligned GZ classes are problem_* or artifact_*. [^3] Another possibility is that the supervised optimization objective (i.e. aligning to GZ) can (counterintuitively) cram class-separating features into a low-dimensional subspace (e.g., Grigg et al., 2021). In other words, the supervised objective can cause embeddings to span a small number of known, dominant morphological classes, whereas the self-supervised case allows representations to disentangle and occupy a broader space of potentially novel features.
@@ 4 Discussion
Limitations. Unfortunately, some SAE features remain difficult to interpret, despite the benefits of using Matryoshka SAEs to combat feature splitting and absorption (Chanin et al., 2024; Bussmann et al., 2025). Still, SAEs surpass our PCA baseline approach. PCA features become unidentifiable after $\sim 30$ features and show no alignment with GZ. In contrast, our SAE delivers an order of magnitude more aligned features (see Appendix A for more details).
Scientific Discovery. SAEs can extract interpretable features from both supervised and self-supervised models at scale (Gao et al., 2024). Sparse activations are steerable and can be readily interpreted through automated means (Cunningham et al., 2023; O’Neill et al., 2024; Ye et al., 2024). Although there are concerns about whether SAEs are effective at learning known concepts, they excel at identifying new features compared to other methods (Muhamed et al., 2024; Movva et al., 2025; Marks et al., 2025; Peng et al., 2025). Thus, SAEs have tremendous potential to aid scientific discovery by categorizing and sifting through next-generation astronomical datasets (see also, e.g., Lochner and Bassett 2021; O’Ryan and Gómez 2025). By extracting features from self-supervised embedding models, SAEs can act as scalable discovery engines that surface rare or anomalous astrophysical phenomena for future targeted investigation.